f(x)=9x+7 and g(x)=5x composite function and state domain. That is what the question says nothing else is given but this .
a. (f o g)(x)
Thank you

Found 2 solutions by Fombitz, fractalier:
Answer by Fombitz(32388)
Answer by fractalier(6550)

(f o g)(x) is read "f composition g of x" but means "f of g of x."
We can write
(f o g)(x) = f(g(x))
so that here
f(g(x)) = f(5x) = 9(5x) + 7 = 45x + 7